OpenStudy (anonymous):

Find the product of (x + 5)^2

OpenStudy (ash2326):

Use the relation \[(a+b)^2=a^2+b^2+2ab\]

OpenStudy (anonymous):

x^2 + 10x + 25

OpenStudy (ash2326):

@ayushprasad Please don't give direct answers. Explain the asker how you got this or give a hint. Thanks

Parth (parthkohli):

Or possibly,\[(a + b)^2 = (a + b)(a + b) = a( a + b) + b(a + b)\]